OpenStudy (anonymous):

(Medal WILL be given!) Rewrite in vertex form y = x^2 + 14x + 50

OpenStudy (tkhunny):

Please don't tag the world. We call that spam. Why don't you "Complete the Square" and show us how it's done?

OpenStudy (anonymous):

i dont know how. thats why im looking for help?

OpenStudy (anonymous):

@tkhunny ^^

OpenStudy (tkhunny):

There's the problem. I don't believe that. If you have no idea whatsoever, you are in the wrong class or you haven't been paying attention. I can't fix that. Completing the Square uses this principle: \((a + b)^{2} = a^{2} + 2ab + b^{2}\) Vertex form looks like this: \(4p(y-k) = (x-h)^{2}\) Those are all the tools you need. Go!